The problem I am trying to figure out is: (a-2)^2=8 I have to use the square root property. I have no idea what I am doing wrong but this is what I have so far.sqrt[(a-2)^2]= +or-sqrt[8]
=a-2=+or-2
==a-2+2=3+or-2
a=3+or-2
a=3+2
a=5
or
a=3-2
a=1
Could you please tell me I have done wrong. And could you show the steps and explain this to me. I just don't understand. Thank you.

Answer by Alan3354(69443)

The problem I am trying to figure out is: (a-2)^2=8 I have to use the square root property. I have no idea what I am doing wrong but this is what I have so far.sqrt[(a-2)^2]= +or-sqrt[8]
=a-2=+or-2
==a-2+2=3+or-2
a=3+or-2
a=3+2
a=5
or
a=3-2
a=1
Could you please tell me I have done wrong. And could you show the steps and explain this to me. I just don't understand. Thank you.
----------------
Take the sqrt of both sides:
(a-2) = +/- sqrt(8) 2 is not the sqrt of 8.
So a = 2 + sqrt(8) and 2 - sqrt(8)
sqrt(8) is 2*sqrt(2), you can sub that in or not.